# Fine-Grained Retrieval-Augmented Generation for Visual Question Answering

### v1

@@ 4.1 Dataset
Table 1: Main results of the experiment. And indicates that the result is from experiments conducted on the full version of the test set, sourced respectively from Hu et al. (2023a) and Chen et al. (2023). Except for the SOTAs, which are trained, other methods are performed in a zero-shot scenario.
Model Dataset
OVENs INFO SEEKs OK-VQA E-VQA
SOTA (Trained) 21.70 22.10 66.10 19.42
LlaVa NEXT-7b 9.51 6.37 73.33 10.51
LlaVa NEXT-7b + KU-RAG 10.80 9.09 73.07 11.00
Llama 3.2-11b 18.50 19.07 68.53 15.04
Llama 3.2-11b + KU-RAG 18.83 19.04 68.89 15.22
GPT-4o 22.30 36.05 75.52 15.17
GPT-4o + KU-RAG 26.50 38.35 77.23 26.16
@@
Table 2 shows some characteristics of each dataset. The more stars in question granularity, the finer the question. The higher the popularity of knowledge, the more general it is, meaning the MLLM is more likely to have learned it during pre-training. Note that since our method is conducted in a zero-shot setting, we only selected the test sets of these datasets. Due to the large size of the original test sets for OVEN and INFO SEEK, we sampled some examples using an arithmetic sequence for testing, and the term ‘$s$’ is used in Table  2 and subsequent experimental results to indicate this.
Table 2: Characteristics of different dataset.
@@ 4.2 Baseline
For the selection of baselines, we chose representative multimodal large language models (MLLMs) with different parameter sizes, including the closed-source model GPT-4o, as well as the open-source models Llava NEXT-7b Liu et al. (2024) and Llama 3.2-11b. Due to variations in the formats and objectives of each dataset, there is no single unified state-of-the-art (SOTA) model across all of them. To ensure a fair comparison, we select the best-performing model for each dataset as its respective SOTA baseline Chen et al. (2022, 2023); Hu et al. (2023a). For detailed information about the baselines, please refer to Appendix B.1.
@@ 4.3 Experimental Setting
Our experiments were conducted on RTX 4090 GPUs. GPT-4o uses the base version of the API interface, while LlaVa and Llama conduct experiments using the Hugging Face transformers library  [^1] https://huggingface.co/docs/transformers/main. In our methods’ settings, OVEN, INFO SEEK, and OK-VQA all use entities as the knowledge unit, while E-VQA uses events as the knowledge unit. For the recall of knowledge units and chunks, the top-k is set to 3. For the experiment evaluation, we used accuracy as the metric.
@@ 4.4 Main Result
As shown in Table  1, we have the following findings.
Zero-shot Capability of MLLMs: MLLM demonstrates remarkable zero-shot capability, especially in image understanding and reasoning. Compared to the previous SOTA model for KB-VQA, GPT-4o shows improvements of 0.6%, 13.95%, and 9.42% on the OVEN, INFO SEEK, and OK-VQA datasets, respectively. This is mainly due to the extensive world knowledge accumulated during MLLM’s pre-training phase. However, since the E-VQA dataset involves less popular news knowledge, MLLM’s performance in this area is not as strong as that of the specially trained SOTA model.
Superior Performance of MLLM+KU-RAG: Our method, MLLM+KU-RAG, performs excellently across all datasets. In a zero-shot scenario, even without reviewing the training set knowledge, GPT-4o+KU-RAG outperforms the existing SOTA models by 4.8%, 16.25%, 11.13%, and 6.74% on the four datasets, respectively, validating the powerful performance of our approach.
Enhancement of MLLM by KU-RAG: Combining KU-RAG with GPT-4o results in performance improvements of 4.2%, 2.3%, 1.7%, and 10.99% on the respective tasks. The largest improvement is seen in the E-VQA dataset, as it involves less popular knowledge, and the new knowledge provided by KU-RAG significantly enhances model performance. In contrast, the improvement on the OK-VQA dataset is smaller because it involves open-domain general knowledge, which MLLM may have already encountered during pre-training, allowing it to answer effectively.
However, for smaller MLLMs like Llava NEXT-7b and Llava 3.2-11b, the improvement from RAG is not very significant, with increases of less than 3% across datasets. This may be due to the limited parameter size, which makes them less capable than GPT-4o in handling the relationship between original and retrieved information.
4.5 Ablation Result
@@
Effectiveness of GPT-4o+KU-RAG: The GPT-4o+KU-RAG method consistently achieves the highest performance, which demonstrates that every part of this method is indispensable.
@@
Impact of Removing KCC: Removing KCC and using single-turn dialogue markedly reduces model performance across four datasets, with decreases of 6%, 18.79%, 8.3%, and 7.97%, respectively. Except for the E-VQA dataset, the model’s performance is inferior to using only GPT-4o. This likely occurs because the model struggles to effectively focus on the original question’s image and manage the logical relationships between the query information, its own knowledge, and the retrieved knowledge. Consequently, some questions, which the model could originally answer correctly, are answered incorrectly due to interference from the injected information.
@@
Impact of Removing KU: Removing KU and adopting a coarse-grained, single-modality retrieval approach results in a slight performance drop across datasets, with the most significant decrease observed in the INFO SEEK dataset (9.04%). This is partly because INFO SEEK requires matching detailed image content and background knowledge, and converting the original image to captions loses a substantial amount of visual information. As illustrated by examples in Figure  1, it’s challenging to accurately match “Karnin Lift Bridge” using just the text “bridge," let alone find corresponding background knowledge. Furthermore, introducing incorrect knowledge adds noise, impeding the MLLM’s reasoning process and leading to erroneous results. The smallest performance drop is observed in the E-VQA dataset (1.65%), likely because, in this dataset, the images primarily serve to supplement information, allowing text-only retrieval to still achieve reasonably good matches.
Figure 4: The Results of Ablation Study
@@ 5 Conclusion
In this paper, we introduce the Knowledge Unit Retrieval-Augmented Generation (KU-RAG) method, aimed at enhancing MLLMs by incorporating fine-grained retrieval of domain-specific knowledge. To improve the effectiveness of retrieval, we propose the concept of “knowledge units”, which allows for more targeted access to relevant information. Furthermore, we design a knowledge correction chain strategy to verify and refine the retrieved knowledge, which can mitigate errors and hallucinations, enhancing the overall reliability and coherence of the generated answers in VQA tasks. Our experimental results demonstrate significant performance gains across multiple KB-VQA benchmarks, highlighting the effectiveness of our approach. Future research directions could explore dynamic knowledge updates stategy to improve multimodal retrieval and semantic integration.
Limitations
Our method exhibits exceptional performance enhancement on large-scale MLLMs such as GPT-4o. However, as evident from the main experimental results, the degree of improvement is comparatively less pronounced when applied to smaller parameter MLLMs. This observation highlights a potential limitation of our approach in effectively scaling down its benefits for smaller models. To address this, future research could focus on refining and optimizing KU-RAG to better accommodate smaller language models, aiming to unlock significant performance gains across a broader range of model sizes.
Appendix A Database Management
Knowledge Addition: When new knowledge is introduced, we directly chunk the corresponding text. The encoded text vectors are then stored in the existing text vector database $D_{c}$. Similarly, for images within the knowledge base, we encode them and store the resulting vectors in the image vector database $D_{i}$.
Knowledge Unit Management: After performing operations of the raw data, it is necessary to consider whether corresponding actions need to be taken for the knowledge unit.
Appendix B Experimental Supplement
Model Dataset
OVENs INFO SEEKs OK-VQA E-VQA
Llama 18.50 19.07 68.53 15.04
Llama + KU-RAG w/o KCC 18.56 18.19 68.25 23.15
Llama + KU-RAG 18.83 19.04 68.89 15.22
LlaVa 9.51 6.37 73.33 10.51
LlaVa + KU-RAG w/o KCC 7.01 5.75 66.35 12.79
LlaVa + KU-RAG 10.08 9.09 73.07 11.00
GPT-4o 22.30 36.05 75.52 15.17
GPT-4o + KU-RAG w/o KCC 20.42 19.56 68.93 18.20
GPT-4o + KU-RAG 26.50 38.35 77.23 26.16
Table 3: Futher Analysis on KCC.
B.1 Baseline
For the MLLMs:
  • GPT-4o: A closed-source MLLM launched by OpenAI, with powerful multimodal content understanding and reasoning capabilities.
  • Llama 3.2-11b: The 11b parameter version of the latest Llama model proposed by Meta AI, also an open-source MLLM.
B.2 Setup and Environment
The experiments were conducted in a zero-shot setting using 4 RTX 4090 GPUs, with PyTorch version 2.3.0. For GPT-4o, we used the interface of the GPT-4o base model. For LLaVa NEXT-7b and Llama 3.2-11b models, we used the Hugging Face Transformers library, version 4.46.3. The LLaVa NEXT-7b model used the weight file ‘llava-v1.6-mistral-7b-hf’, while the Llama 3.2-11b model loaded the weight file ‘Llama-3.2-11B-Vision-Instruct’.
B.3 Futher Analysis on KCC
We conducted a further analysis on the impact of removing the Knowledge Correction Chain (KCC), as shown in Table 3. Based on the results, for the OVEN, INFO SEEK, and OK-VQA datasets, the performance of the models, whether it’s the smaller models like Llama and LlaVa, or the powerful large language model GPT-4o, all decreased to varying degrees after removing KCC. This suggests that KCC plays a significant role in enhancing model performance, especially for GPT-4o, where the performance drop after removing KCC is the most substantial, indicating that the gain from KCC is strongest for this model.
Additionally, after removing KCC, the models are influenced by the external knowledge retrieved, which leads them to disregard their own internal knowledge. For the OVEN, INFO SEEK, and OK-VQA datasets, the performance without KCC did not surpass the performance of the models used individually. This indicates that KCC helps balance the relationship between the retrieved knowledge and the model’s own knowledge, thereby improving overall performance.
It’s worth noting that for the E-VQA dataset, Llama and LlaVa experienced a performance drop when KCC was introduced. The main reason for this is that the knowledge in the E-VQA dataset is more specific and less general, so the answers to the questions mainly rely on the knowledge retrieved. However, due to the smaller model sizes of Llama and LlaVa, they struggle to effectively manage the conflict between their internal knowledge and the retrieved knowledge, leading to a decline in performance. In contrast, GPT-4o, with its larger model capacity, is able to handle these conflicts more effectively, resulting in better performance on the E-VQA dataset.

### v2

@@ 6.1 Dataset
Table 1 shows some characteristics of each dataset. The more stars in question granularity, the finer the question. The higher the popularity of knowledge, the more general it is, meaning the MLLM is more likely to have learned it during pre-training. Note that since our method is conducted in a zero-shot setting, we only selected the test sets of these datasets. Due to the large size of the original test sets for OVEN and INFO SEEK, we sampled some examples using an arithmetic sequence for testing, and the term ‘$s$’ is used in Table  1 and subsequent experimental results to indicate this.
Table 1: Characteristics of Different Dataset
@@
Table 2: Main results of the experiment. And indicates that the result is from experiments conducted on the full version of the test set, sourced respectively from Hu et al. (2023a) and Chen et al. (2023). Except for the SOTAs, which are trained ([Uncaptioned image]), other methods are performed in a zero-shot scenario ([Uncaptioned image]).
Model Dataset
OVENs INFO SEEKs OK-VQA E-VQA
SOTA[Uncaptioned image] 21.70 22.10 66.10 19.42
LlaVa NEXT-7b[Uncaptioned image] 9.51 6.37 73.33 10.51
LlaVa NEXT-7b + KU-RAG[Uncaptioned image] 10.80 9.09 73.07 11.00
Qwen 2.5-VL-32b[Uncaptioned image] 28.44 27.01 73.04 14.82
Qwen 2.5-VL-32b + KU-RAG[Uncaptioned image] 30.42 27.03 73.46 21.10
GPT-4o[Uncaptioned image] 22.30 36.05 75.52 15.17
GPT-4o + KU-RAG[Uncaptioned image] 26.50 38.35 77.23 26.16
@@ 6.2 Baseline
For the MLLMs:
  • GPT-4o: A closed-source MLLM launched by OpenAI with powerful multimodal content understanding and reasoning capabilities.
  • Qwen 2.5-VL-32b: A 32b-parameter open-source MLLM developed by Alibaba, capable of advanced vision-language understanding and reasoning.
@@ 6.3 Experimental Setting
Our experiments were conducted on RTX 4090 GPUs. GPT-4o and Qwen 2.5-VL-32b use the base version of the API interface, while LlaVa conducts experiments using the Hugging Face transformers library  [^1] https://huggingface.co/docs/transformers/main. The LLaVa NEXT-7b model used the weight file ‘llava-v1.6-mistral-7b-hf’. In our method’s settings, OVEN, INFO SEEK, and OK-VQA all use entities as the knowledge unit, while E-VQA uses events as the knowledge unit. For the recall of knowledge units and chunks, the top-k is set to 3. For the experiment evaluation, we used accuracy as the metric.
@@ 6.4 Main Result
As shown in Table  2, we have the following findings.
Zero-shot Capability of MLLMs. MLLMs exhibit remarkable zero-shot capabilities in image understanding and reasoning. Compared to the previous SOTA model for KB-VQA, GPT-4o shows improvements of 0.6%, 13.95%, and 9.42% on the OVEN, INFO SEEK, and OK-VQA datasets, respectively, benefiting from its extensive world knowledge acquired during pre-training. However, as the E-VQA dataset involves less commonly known news-related knowledge, GPT-4o struggles to outperform the trained SOTA model in this scenario.
Smaller models, such as LLaVA NEXT-7b and Qwen 2.5-VL-32b, also demonstrate strong zero-shot reasoning ability but fall short of GPT-4o. Notably, Qwen 2.5-VL-32b performs significantly better than LLaVA NEXT-7b across all datasets, highlighting its stronger multi-modal understanding.
Superior Performance of MLLM+KU-RAG. Our proposed method, MLLM+KU-RAG, achieves superior results across all datasets. In a zero-shot scenario, even without prior exposure to the training set knowledge, GPT-4o+KU-RAG outperforms the existing SOTA models by 4.8%, 16.25%, 11.13%, and 6.74% on the four datasets, respectively, validating its strong reasoning and retrieval augmentation capabilities.
Furthermore, Qwen 2.5-VL-32b+KU-RAG achieves the highest zero-shot accuracy among mid-sized models, surpassing LLaVA NEXT-7b+KU-RAG by a large margin. This suggests that models with better intrinsic multi-modal alignment can benefit more from knowledge retrieval.
Enhancement of MLLM by KU-RAG. Integrating KU-RAG with GPT-4o leads to performance gains of 4.2%, 2.3%, 1.7%, and 10.99% across the respective datasets. The most significant improvement is observed on E-VQA, where KU-RAG provides crucial missing knowledge, while the gain on OK-VQA is smaller due to the model’s inherent knowledge coverage.
For smaller models, the benefits of KU-RAG are also evident, though less pronounced than in GPT-4o. Qwen 2.5-VL-32b sees improvements of 1.98%, 2.1%, 0.4%, and 6.28% across the four datasets, showing its ability to integrate retrieved information effectively. In contrast, LLaVA NEXT-7b benefits the least from RAG, suggesting that both model architecture and parameter scale play key roles in utilizing external knowledge efficiently.
These results highlight that KU-RAG is particularly effective when paired with strong multi-modal models, significantly enhancing factual consistency and knowledge coverage in zero-shot settings.
Figure 5: The Results of Ablation Study
6.5 Ablation Study
@@
Effectiveness of GPT-4o+KU-RAG. The GPT-4o+KU-RAG method consistently achieves the highest performance across all four benchmarks, indicating the effectiveness and complementarity of its components. This result highlights the importance of integrating both KU and KCC in the retrieval-augmented generation pipeline.
@@
Impact of Removing KCC. Removing KCC and using single-turn dialogue markedly reduces model performance across four datasets, with decreases of 6%, 18.79%, 8.3%, and 7.97%, respectively. Except for the E-VQA dataset, the model’s performance is inferior to using only GPT-4o. This likely occurs because the model struggles to effectively focus on the original question’s image and manage the logical relationships between the query information, its own knowledge, and the retrieved knowledge. Consequently, some questions that the model could originally answer correctly are answered incorrectly due to interference from the injected information.
@@
Impact of Removing KU. Removing KU and adopting a coarse-grained, single-modality retrieval approach results in a slight performance drop across datasets, with the most significant decrease observed in the INFO SEEK dataset (9.04%). This is partly because INFO SEEK requires matching detailed image content and background knowledge, and converting the original image to captions loses a substantial amount of visual information. As illustrated by examples in Figure  1, it’s challenging to accurately match “Karnin Lift Bridge” using just the text “bridge," let alone find corresponding background knowledge. Furthermore, introducing incorrect knowledge adds noise, impeding the MLLM’s reasoning process and leading to erroneous results. The smallest performance drop is observed in the E-VQA dataset (1.65%), likely because, in this dataset, the images primarily serve to supplement information, allowing text-only retrieval to still achieve reasonably good matches.
Comparison with GPT-4o Only. Notably, the performance of certain ablation variants (particularly the version excluding KCC) underperforms the standalone GPT-4o model on datasets such as INFO SEEK and OK-VQA. This observation reinforces the critical importance of knowledge integration mechanisms: when retrieved content lacks proper orchestration through strategies like KCC or KU, the model may assimilate irrelevant or contradictory information that disrupts its reasoning processes. Our findings demonstrate that effective retrieval-augmented generation depends not only on external knowledge access but more fundamentally on systematic integration frameworks. Specifically, successful implementations require architectures that enable contextual coherence and logical synthesis of retrieved information during the reasoning phase, highlighting the necessity of deliberate fusion strategies over mere knowledge injection.
7 Related Work
7.1 Knowledge-based Visual Question Answering
Knowledge-based Visual Question Answering (KB-VQA) aims to leverage external knowledge to assist in answering questions about images  (Marino et al., 2019; Gardères et al., 2020). In early KB-VQA approaches (Zhu et al., 2020; Gao et al., 2022; Lin and Byrne, 2022), Wikipedia was often used as the external knowledge source for KB-VQA Caffagni et al. (2024), leading to the common adoption of a retriever-reader framework. This framework first retrieves textual knowledge relevant to the question and image, and then “reads” the text to predict the answer. However, this passage dense retrieval method is a unimodal, coarse-grained text-to-text retrieval process, which struggles with specific, fine-grained questions such as visual entities (Hu et al., 2023a), events (Yang et al., 2023), and visual information-seeking questions (Chen et al., 2023).
Since the emergence of LLMs, some methods have explored using implicit knowledge from LLMs in addition to retrieving information from databases like Wikipedia Caffagni et al. (2024). Typically, they convert images into tags or captions and then use GPT to retrieve related knowledge (Gui et al., 2021). However, there is a gap between the query and the LLM’s knowledge source. To address this,  Hu et al. (2023b) proposed a prompt-guided image captioning method that controls the visual entities in generated captions based on textual queries, replacing general captions with question-dependent ones. Although some methods attempt to mitigate the loss of visual information by incorporating visual features (Salaberria et al., 2023) or enriching prompts with candidate answers (Shao et al., 2023), they have not completely solved the issue of information loss. Recently, Hao et al. (2024) introduce a self-bootstrapped visual-language model that refines retrieved knowledge using a selector-answerer framework, significantly improving knowledge selection and QA accuracy, but their method still requires complex training.
Our approach shifts the focus from retrieval optimization to knowledge representation by introducing Knowledge Units, which serve as structured bridges between queries and multimodal knowledge. Rather than treating retrieval as an isolated process, we integrate multimodal information into a unified retrieval and reasoning framework.
7.2 Multimodal Retrieval-augmented Generation
Although LLMs possess strong general knowledge answering capabilities, they still face limitations when dealing with domain-specific knowledge, outdated information, and avoiding hallucinations  (Gao et al., 2023). To address these issues, Retrieval-Augmented Generation (RAG) was developed. RAG enhances the answering ability of LLMs by retrieving relevant document fragments from external knowledge bases. Specifically, the RAG approach involves multiple modules such as data storage, query optimization, document retrieval, and answer generation. The basic process matches user queries with documents from a large external knowledge base, retrieves relevant document fragments, and generates answers by integrating this information through a generation model. This process is similar to the knowledge retrieval mechanisms used in KB-VQA.
Building on this, RAG has evolved to optimize the retrieval and generation processes. For example, GRAG (Graph Retrieval-Augmented Generation) improves the relevance of information and generation quality by emphasizing the importance of subgraph  (Hu et al., 2024). The FiD-RAG (Fusion-in-Decoder RAG) model parallelly fuses multiple retrieved documents during the generation stage, allowing the model to comprehensively integrate background knowledge from different sources  (Izacard and Grave, 2020). Moreover, DPR-RAG (Dense Passage Retrieval RAG) introduces dense retrieval techniques that significantly improve retrieval accuracy, quickly locating highly relevant fragments from large document collections  (Karpukhin et al., 2020).
Unlike these approaches, which primarily refine retrieval mechanisms, our method focuses on the structured representation of multimodal knowledge. By organizing knowledge into Knowledge Units, we establish a persistent, query-aware knowledge structure, ensuring fine-grained, contextually relevant retrieval.
@@ 8 Conclusion
In this paper, we introduce the KU-RAG, aimed at enhancing MLLMs by incorporating fine-grained retrieval of domain-specific knowledge. To improve the effectiveness of retrieval, we propose the concept of “knowledge units”, which allows for more targeted access to relevant information. Furthermore, we design a knowledge correction chain strategy to verify and refine the retrieved knowledge, which can mitigate errors and hallucinations, enhancing the overall reliability and coherence of the generated answers in VQA tasks. Our experimental results demonstrate significant performance gains across multiple KB-VQA benchmarks, highlighting the effectiveness of our approach.
Future research directions may explore dynamic knowledge updates to improve adaptability, and integrate user feedback to enhance retrieval relevance and answer accuracy. Extending KU-RAG to other multimodal tasks like visual dialogue could further demonstrate its generalizability.